We got visitors – University of Copenhagen

NEEM > Field diaries folder > UK diaries 2009 > We got visitors

22 June 2009

We got visitors

Skier 02 in the sky overhead

Skier 02 in the sky overhead.

The 109th had asked us if we would host a visit from a U.S. press delegation, and of course we said yes. As we had complete overcast, we had to launch a weather balloon to determine the cloud base. The cloud base was measured to be more than 3,000 feet, so the plane took off from Kangerlussuaq at 9.40. At 12.38 the plane landed in camp, and we gave the visitors a tour of the activities. The visitors happened to be in the drill trench as an ice core was pulled up. At 14.10 the plane left camp with no difficulty at all. 30 minutes later, it began to snow intensely, reducing visibility to few hundred meters. We had been lucky again.

What we have done today:

  1. Drilling with the NEEM long drill.
  2. Logging is on pause. Last bag before brittle zone: 1021. Depth: 561.55 m.
  3. Processing 21.45 m. Processed depth: 350.90 m.
  4. Measurements in the CFA laboratory. Measurements now cover from 7 m to 75.90 m.
  5. Receiving Skier 02 with a group of press and outreach people.
  6. Unpacked food pallet.
  7. Groomed away holes in skiway with tiller.
  8. Cooling tunnel activated in science trench. Within two hours, temperature dropped from -15°C to -18°C.

Ad.1: This is how the drillers report looks today:
"Stable drilling. Only 2.2 m runs were allowed due to length of available troughs. Tomorrow most of the drilling team, Henry, HC, Seb, Nobby and Sigfus, will fly back. We wish the ones left behind, JiWoong and Alex and the newcomers, Phillipe, Olivier, Frank, Fernando and Mads all possible luck with our easy going drill. The rollers in the pump should be replaced tomorrow. Inclination has been 1 deg or less for several days. We drilled 17.89 m in 6 runs; drillers depth 566.16 m."

Ad.2: The loggers are now taking the fresh cores directly to the buffer to allow for the core to relax for 5 days before it is being logged and cut into sections. The cores from bag 1022 until the brittle zone ends will not be processed in 2009.

Ad.5: The 109th paid us a visit today. Along with press and outreach officers, they brought TV crews and journalists. They also brought our new cook, Brandon, who takes over from Sarah tomorrow and Sune, our coming field office manager in Kangerlussuaq, who got the opportunity to spend a night in camp. As cargo they brought some fuel, some food and ice core boxes. Finally, the crew evaluated the skiway, and it was certified for bad weather approaches. Thank you 109th!. We are now 34 in camp.

Ad.8: The cooling tunnel is already doing its work. The next crew in the science trench will feel the cold; but so will the ice cores, and they benefit.

Weather: Interesting. Completely overcast all day until 23:00, then suddenly clear. Temp. last night -12°C to -10°C by day. At 23:00 a sudden drop to -17°C, 5-10 knots from S and SW. Fromt 17:00 to 20:00 a three hour blow with 15-20 knots. Visibility: 3-5 miles most of the day, reduced to less than one mile during intense snow fall starting 14:40. After 23:00 unrestricted.

Field leader J.P. Steffensen

← Previous entry   Next entry →